# PPFL-RDSN: Privacy-Preserving Federated Learning-based Residual Dense Spatial Networks for Encrypted Lossy Image Reconstruction

### v1

@@ Abstract
Reconstructing high-quality images from low-resolution inputs using Residual Dense Spatial Networks (RDSNs) is crucial yet challenging, particularly in collaborative scenarios where centralized training poses significant privacy risks, including data leakage and inference attacks, as well as high computational costs. We propose a novel Privacy-Preserving Federated Learning-based RDSN (PPFL-RDSN) framework specifically tailored for lossy image reconstruction. PPFL-RDSN integrates Federated Learning (FL), local differential privacy, and robust model watermarking techniques, ensuring data remains secure on local devices, safeguarding sensitive information, and maintaining model authenticity without revealing underlying data. Empirical evaluations show that PPFL-RDSN achieves comparable performance to the state-of-the-art centralized methods while reducing computational burdens, and effectively mitigates security and privacy vulnerabilities, making it a practical solution for secure and privacy-preserving collaborative computer vision applications.
@@
Privacy-Preserving Federated Learning, Differential Privacy, Model Watermarking, Lossy Image Reconstruction.
1 Introduction
@@
Encrypted lossy image reconstruction using neural networks has emerged as a critical area of research to address the need for high-quality image reconstruction while ensuring security, high accuracy, and improved performance [1, 2, 3]. In existing centralized RDSN approaches [4, 5, 6], all the training data is expected to be available in the central server before training the model. Thus, centralized training of machine learning (ML) models requires transferring massive amounts of data from participants or data providers to a central server, which creates significant communication overhead, and increases the attack surface for security and privacy compromises [7]. This creates substantial scaling challenges in scenarios where there is a large pool of participants/data providers. Furthermore, centralizing data from diverse sources increases the risk of data exposure or leakage, privacy breaches, and malicious or adversarial attacks. To address privacy and concerns and scaling challenges in collaborative ML settings, federated learning (FL) was introduced by McMahan et al. [8]. In FL, each participating client independently trains a local ML model using its dataset, thus preventing raw data from being exposed to a central server. After local training, clients send their model updates to an aggregator, which iteratively combines these updates to create a global model. To further protect privacy and maintain the confidentiality of these model updates, various secure and privacy-preserving techniques have been proposed, including Differential Privacy [9, 10, 11, 12], Homomorphic Encryption [13, 14, 15, 16], and Secure Multi-Party Computation [17, 18, 19]. However, implementing these methods introduces additional computational overhead, posing a crucial challenge in balancing privacy guarantees and model accuracy, as well as practically acceptable performance.
@@
Encryption-then-Lossy-Compression (EtLC) schemes [20, 21, 22] have been proposed to securely reconstruct images which typically involve two parties: a sender and a receiver. In these approaches, the receiver must either obtain the Residual Dense Spatial Network (RDSN) model from a third-party cloud server or train it independently; however, both approaches have significant drawbacks. Acquiring the model from an untrusted cloud server allows for potential model poisoning attacks, leading to unreliable/corrupted reconstructions at the receiver side. Conversely, local training of the RDSN imposes substantial computational and communication costs. In addition to these practical issues, encrypted image reconstruction tasks face broader challenges of ensuring robust privacy and security protections, especially in scenarios involving multiple distributed data providers, and minimizing the computational burden on individual clients without compromising model performance. Existing RDSN techniques do not address these privacy and security concerns, leaving sensitive information vulnerable during both model training and inference.
@@
To address these critical issues, this paper proposes a novel Privacy-Preserving Federated Learning-based RDSN (PPFL-RDSN) scheme tailored for encrypted lossy image reconstruction in a collaborative, multi-party environment. PPFL-RDSN introduces a federated architecture comprising a central server, responsible for aggregating model updates, and multiple clients collaboratively training an enhanced RDSN model through locally prepared tunable parameters. Building upon and customizing the basic RDSN architecture from Wang et al. [20], our approach enables effective distributed training, achieves high image reconstruction quality while ensuring privacy protection. The proposed PPFL-RDSN framework extends the conventional encryption-then-lossy-compression (EtLC) [23] paradigm from the traditional two-party setup to a multi-party collaborative setting. Within this extended framework, each client dynamically assumes the roles of sender or receiver, enabling distributed high-resolution image reconstruction without collecting sensitive data from individual providers/participants in a central location. Given that direct integration of differential privacy (DP) noise severely degrades reconstruction quality, our approach leverages Local Differential Privacy (LDP), strategically injecting Gaussian noise into selected feature spaces or parameter updates. This ensures robust privacy without compromising reconstruction quality. Additionally, the proposed mechanism incorporates a watermarking verification mechanism to safeguard model integrity during aggregation.
@@
  • We propose PPFL-RDSN, a novel federated learning framework that extends encrypted lossy image reconstruction from two-party to multi-party collaboration, emphasizing robust privacy preservation.
  • We introduce an enhanced RDSN architecture customized for federated learning, significantly improving image reconstruction quality and reducing training cost.
  • We develop a strategic Local Differential Privacy approach specifically designed for encrypted image reconstruction, balancing strong privacy protections with minimal accuracy loss.
  • We embed a watermarking-based integrity verification mechanism, ensuring clients can reliably verify the authenticity and integrity of aggregated global models.
The experimental results show that our approach achieves an accuracy comparable (as measured by PSNR/SSIM metrics) to state-of-the-art methods while improving training efficiency by approximately 6.56%. More importantly, our framework provides mitigation against inference attacks and model poisoning or alteration attempts, ensuring a balance between utility and privacy in real-world scenarios.
2 Background
In this section, we overview the key concepts and introduce the notation used in our proposed PPFL-RDSN framework.
2.1 Deep learning-based RDSN
The Residual Dense Spatial Network (RDSN), proposed by Wang et al. [20], is a specialized deep-learning architecture designed explicitly for efficiently reconstructing encrypted-then-lossy-compressed images [24, 25, 26]. RDSN extends the Residual Dense Network (RDN) family [27] by blending two complementary mechanisms: 1). Residual learning, which predicts the high‑frequency residual between the degraded input and its pristine counterpart, enabling efficient flow of low‑frequency information through skip connections; and 2). Dense connectivity, which concatenates feature maps from all preceding layers, encourages feature reuse and alleviates vanishing‑gradient issues. Within each RDN block, the residual path expresses the network’s output as:
@@ 2.1 Deep learning-based RDSN
where $\mathbf{x}$ is the input, and $\mathcal{F}$ encapsulates convolutional, activation, and normalization operations parameterized by weights $\left\{W_{i}\right\}$. Each dense block enhances feature propagation by concatenating the outputs of all preceding layers, mathematically expressed as:
@@
where $H_{l}$ denotes the transformation at layer $l$. Further, the RDSN model specifically incorporates three crucial enhancements: an improved U-net architecture [28], a Global Skip Connection (GSC), and Uniformly Downsampling Constraints (UDC) [29]. The refined U-net employs a sub-pixel convolution operation, defined by: $\mathbf{Y}=\operatorname{PS}(\mathbf{W}*\mathbf{I})$, where $\mathbf{W}$ represents the convolutional kernel, * denotes convolution, and $\operatorname{PS}(\cdot)$ performs the pixel shuffle operation that rearranges elements to enhance spatial resolution. The Global Skip Connection significantly enhances model efficiency by directly propagating essential features from input to deeper network layers, preserving crucial structural information. This relationship is described as:
@@
Collectively, these components enable RDSN to achieve robust and computationally efficient reconstruction performance in high-resolution imaging tasks.
2.2 Federated Learning
McMahan et al. propose Federated Learning [8] that allows multiple clients to collaboratively train a global model by combining the local models of these clients trained on their local data (which do not need to be shared with a central server), and an aggregator combining these local models to build the global model. Their method for federated learning of deep networks uses iterative model averaging and achieves a 10-100x reduction in communication rounds compared to the synchronized stochastic gradient descent approach [30, 31]. Each client or device $k$ updates its local model parameters as follows:
$w_{k}^{(t+1)}=w_{k}^{(t)}-\eta\nabla F_{k}(w_{k}^{(t)},\mathcal{B}_{k})$
where: $w_{k}^{(t)}$ are the model parameters of device $k$ at iteration $t$, $\eta$ is the learning rate $\nabla,F_{k}(w_{k}^{(t)},\mathcal{B}_{k})$ is the gradient computed on the local batch of data $\mathcal{B}_{k}$ When all clients upload their local models each round (i.e., share with the aggregator), through Fed-Aggregation, the central server or aggregator aggregates all the local models for each iteration:
$w^{(t+1)}=\sum_{k=1}^{K}\frac{n_{k}}{n}w_{k}^{(t+1)}$
where: $w^{(t+1)}$ is the global model parameters at iteration $t+1$, $K$ is the total number of participating devices, $n_{k}$ is the number of data points on device $k$, and $n$ is the total number of data points across all devices, i.e., $n=\sum_{k=1}^{K}n_{k}$. After aggregation, the central server distributes the latest model updates to each client. This approach [32, 33] aims to enhance data privacy and reduce communication overhead by only sharing model updates instead of raw data. Federated Learning [34, 35, 36, 37, 38, 39, 40, 41, 42] can be used in various distributed application domains, such as healthcare, image processing, large language models, and explainable AI.
@@ 2.2 Federated Learning
3 Proposed PPFL-RDSN Framework
In this section, we provide an overview of the PPFL-RDSN framework in Figure 1, with key components as follows:
  • Local Training: Within the PPFL-RDSN framework, each client operates its local training system, maintaining exclusive access to its dataset without sharing data with other clients. During local training, Gaussian noise is strategically applied to the parameter updates to balance data utility and privacy, mitigating risks of data leakage. Additionally, each client embeds a watermark into its local RDSN model during training for subsequent integrity verification.
  • Model Aggregation and Distribution: After training, clients securely transmit their local models to a central server equipped with a secure aggregator. The server performs model aggregation using Federated Averaging (FedAvg) [43], combining local models into a global model based on weighted contributions.
  • Local EtLC: Once aggregated, the global model is distributed back to each client. Upon receipt, clients execute a model integrity check using their embedded watermark. Successful verification leads to performance testing, involving inputting lossy or low-resolution images into the global model. Depending on the test outcomes, the client either retains the global model or reinitiates local training iterations. Additionally, the Local EtLC subsystem allows clients to securely exchange encrypted and compressed low-resolution images, but how to share secret keys is out of scope this time. Upon request, these images are decrypted, decompressed, and reconstructed into higher-resolution outputs using the verified global RDSN model. This process ensures collaborative enhancement while maintaining data security and integrity throughout the federated learning network.
3.1 Threat Modeling and Mitigation
We analyze two classes of adversaries: 1): an untrusted aggregator and 2): honest‑but‑curious clients. We show how every identified threat is countered by a specific PPFL‑RDSN component.
Privacy Threats: Membership, inversion, and property inference attacks exploit shared gradients to uncover private training data [44, 45, 46]. These threats arise from both untrusted aggregators and curious clients who might use the global model parameters to uncover private details.
Integrity / Security Threats: Poisoning attacks [47] use malicious gradient updates that can stealthily degrade or subvert the global model; Model substitution or parameter tampering attacks [48, 49] utilize the untrusted server, which may replace the global model with a Trojan version.

### v3

@@ Abstract
Reconstructing high-quality images from low-resolution inputs using Residual Dense Spatial Networks (RDSNs) is crucial yet challenging. It is even more challenging in centralized training where multiple collaborating parties are involved, as it poses significant privacy risks, including data leakage and inference attacks, as well as high computational and communication costs. We propose a novel Privacy-Preserving Federated Learning-based RDSN (PPFL-RDSN) framework specifically tailored for encrypted lossy image reconstruction. PPFL-RDSN integrates Federated Learning (FL), local differential privacy, and robust model watermarking techniques to ensure that data remains secure on local clients/devices, safeguards privacy-sensitive information, and maintains model authenticity without revealing underlying data. Empirical evaluations show that PPFL-RDSN achieves comparable performance to the state-of-the-art centralized methods while reducing computational burdens, and effectively mitigates security and privacy vulnerabilities, making it a practical solution for secure and privacy-preserving collaborative computer vision applications.
@@
Privacy-Preserving Federated Learning, Differential Privacy, Watermarking, Lossy Image Reconstruction
I Introduction
@@
Encrypted lossy image reconstruction using neural networks has emerged as a critical area of research to address the need for high-quality image reconstruction while ensuring security, high accuracy, and improved performance [1, 2, 3]. It involves an image processing pipeline where images are protected before or during compression and later reconstructed from the encrypted, compressed representation by learned models. The aim is to keep content confidential in storage and transit while retaining bitrate savings and the quality gains of neural post-processing, which is valuable for privacy-sensitive, bandwidth-limited application settings such as telemedicine, remote sensing, and cloud photo services. In existing centralized RDSN approaches [4, 5, 6], all the training data is expected to be available in the central server before training the model. Thus, centralized training of machine learning (ML) models requires transferring massive amounts of data from participants or data providers to a central server, which creates significant communication overhead, and increases the attack surface for security and privacy compromises [7]. This creates substantial scaling challenges in scenarios where there is a large pool of participants/data providers. Furthermore, centralizing data from diverse sources increases the risk of data exposure or leakage, privacy breaches, and malicious or adversarial attacks. To address privacy and concerns and scaling challenges in collaborative ML settings, federated learning (FL) was introduced by McMahan et al. [8]. In FL, each participating client independently trains a local ML model using its dataset, thus preventing raw data from being exposed to a central server. After local training, clients send their model updates to an aggregator, which iteratively combines these updates to create a global model. To further protect privacy and maintain the confidentiality of these model updates, various secure and privacy-preserving techniques have been proposed, including Differential Privacy [9, 10, 11, 12], Homomorphic Encryption [13, 14, 15, 16], and Secure Multi-Party Computation [17, 18, 19]. However, implementing these methods introduces additional computational overhead, posing a crucial challenge in balancing privacy guarantees and model accuracy, as well as practically acceptable performance.
@@
Encryption-then-Lossy-Compression (EtLC) schemes [20, 21, 22] have been proposed to securely reconstruct images where two parties are typically involved: a sender and a receiver. EtLC schemes first encrypt an image first and then apply a standard or learned compressor. This ensures that untrusted networks handle only encrypted payloads while staying compatible with off-the-shelf codecs and delivery systems. In practice, EtLC enables secure sharing without sacrificing compression efficiency and pairs naturally with learned reconstructor networks. In these approaches, the receiver must either obtain the Residual Dense Spatial Network (RDSN) model from a third-party, such as a cloud server, or train it independently; however, both approaches have significant drawbacks. Acquiring the model from an untrusted cloud server allows for potential model poisoning attacks, leading to unreliable/corrupted reconstructions at the receiver side. Conversely, local training of the RDSN imposes substantial computational and communication costs. In addition to these practical issues, encrypted image reconstruction tasks face broader challenges of ensuring robust privacy protection and security, especially in scenarios involving multiple distributed data providers, and minimizing the computational burden on individual clients without compromising model performance. Existing RDSN techniques do not address these privacy and security concerns, leaving sensitive information vulnerable during both model training and inference.
@@
To address these critical issues, we propose a novel Privacy-Preserving Federated Learning-based RDSN (PPFL-RDSN) scheme tailored for encrypted lossy image reconstruction in a collaborative, multi-party environment. PPFL-RDSN introduces a federated architecture comprising a central server, responsible for aggregating model updates, and multiple clients collaboratively training an enhanced RDSN model through locally controlled tunable parameters. Building upon and customizing the basic RDSN architecture from Wang et al. [20], our approach enables effective distributed training, achieves high image reconstruction quality while ensuring privacy protection. The proposed PPFL-RDSN framework extends the conventional encryption-then-lossy-compression (EtLC) [23] paradigm from the traditional two-party setup to a multi-party collaborative setting. Within this extended framework, each client can dynamically assume the roles of sender or receiver, enabling distributed high-resolution image reconstruction without collecting sensitive data from individual providers/participants in a central location. Given that direct integration of differential privacy (DP) noise severely degrades reconstruction quality, our approach leverages Local Differential Privacy (LDP), strategically injecting Gaussian noise into selected feature spaces or parameter updates. This ensures robust privacy without compromising reconstruction quality. Additionally, the proposed mechanism incorporates a watermarking verification mechanism to safeguard model integrity during aggregation.
@@
  • We propose PPFL-RDSN, a novel federated learning framework that extends encrypted lossy image reconstruction from two-party to multi-party collaboration, providing robust privacy protection.
  • We introduce an enhanced RDSN architecture tailored for federated learning, which significantly improves image reconstruction quality and reduces training costs.
  • We develop and employ a Local Differential Privacy approach specifically designed for encrypted image reconstruction, that balances strong privacy protections with minimal accuracy loss.
  • We integrate a watermarking-based integrity verification mechanism to ensure that clients can reliably verify the authenticity and integrity of aggregated global models.
The experimental results show that our approach achieves an accuracy comparable (as measured by PSNR/SSIM metrics) to state-of-the-art methods while improving training efficiency by approximately 6.56%. More importantly, our framework provides mitigation against inference attacks and model poisoning or alteration attacks, ensuring a balance between utility and privacy in real-world scenarios.
II Background and Related Work
In this section, we overview the key concepts and notation underlying our PPFL-RDSN framework for reconstructing encrypted-then-lossy-compressed images. We first summarize the Residual Dense Spatial Network (RDSN) architecture used for high-fidelity reconstruction, and then revisit federated learning (FL) with a focus on the privacy and integrity mechanisms—most notably differential privacy (DP) and model verification—that are essential when training complex image models across heterogeneous clients.
II-A Deep learning-based RDSN
The Residual Dense Spatial Network (RDSN), proposed by Wang et al. [20], is a specialized deep-learning architecture designed explicitly for efficiently reconstructing encrypted-then-lossy-compressed images [24, 25, 26]. RDSN extends the Residual Dense Network (RDN) family [27] by blending two complementary mechanisms: (1) Residual learning, which predicts the high-frequency residual between the degraded input and its original counterpart, enabling efficient flow of low-frequency information through skip connections; and (2) Dense connectivity, which concatenates feature maps from all preceding layers, encouraging feature reuse and alleviating vanishing-gradient issues. Within each RDN block, the residual path expresses the network’s output as [??]:
@@ II-A Deep learning-based RDSN
where $\mathbf{x}$ is the input, and $\mathcal{F}$ encapsulates convolutional, activation, and normalization operations parameterized by weights $\left\{W_{i}\right\}$. Each dense block enhances feature propagation by concatenating the outputs of all preceding layers, as follows:
@@
where $H_{l}$ denotes the transformation at layer $l$.
Further, the RDSN model specifically incorporates three crucial enhancements: an improved U-Net architecture [28], a Global Skip Connection (GSC), and Uniformly Downsampling Constraints (UDC) [29]. The refined U-Net employs a sub-pixel convolution operation, defined by $\mathbf{Y}=\operatorname{PS}(\mathbf{W}*\mathbf{I})$, where $\mathbf{W}$ represents the convolutional kernel, * denotes convolution, and $\operatorname{PS}(\cdot)$ performs the pixel shuffle operation that rearranges elements to enhance spatial resolution. The Global Skip Connection significantly improves efficiency by directly propagating essential features from input to deeper layers, preserving crucial structural information:
@@
While encryption-then-lossy-compression pipelines have been explored to prevent direct content inspection [23, 30], these prior solutions typically assume a two-party setting with a trusted receiver. Scaling to multi-participant collaboration introduces coordination and trust challenges around both the encrypted data and the global model’s authenticity. The closest existing work related to our proposed approach is the RDSN approach proposed in [20], which targets reconstruction quality but does not integrate federated contrastive training [31], enhanced privacy protections, or cryptographic verification. Our framework operates in this distributed, multi-party setting, where an acceptable, higher reconstruction quality must be preserved while providing privacy and integrity guarantees.
II-B Federated Learning
Federated Learning (FL) enables multiple clients to collaboratively train a global model by locally updating on their private data and sharing updates with a central aggregator, thereby reducing raw-data exposure [8]. Concretely, each client or device $k$ performs local optimization
$w_{k}^{(t+1)}=w_{k}^{(t)}-\eta\nabla F_{k}\!\big(w_{k}^{(t)},\mathcal{B}_{k}\big),$
where $w_{k}^{(t)}$ are the model parameters of device $k$ at iteration $t$, $\eta$ is the learning rate, and $\nabla F_{k}(\cdot,\mathcal{B}_{k})$ is the gradient computed on local mini-batch $\mathcal{B}_{k}$. The aggregator then forms the new global model via weighted averaging (FedAvg [32]):
$w^{(t+1)}=\sum_{k=1}^{K}\frac{n_{k}}{n}\,w_{k}^{(t+1)},$
where $K$ is the number of participating devices, $n_{k}$ is the number of data points on device $k$, and $n=\sum_{k=1}^{K}n_{k}$. This iterative averaging reduces the number of communication rounds relative to synchronized SGD [33, 34], and has been deployed across diverse domains such as image processing, large language models, and explainable AI—where data locality and heterogeneity are the norm [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
However, two considerations become critical for high-fidelity image reconstruction in FL. First, most of the existing FL work for image analysis has focused on comparatively simpler objectives such as classification or segmentation [45, 46, 47]. Reconstruction demands preserving subtle textures and high-frequency details under distribution shift and client heterogeneity, which stresses both the optimization and the communication related to computational costs. Second, although FL reduces direct data sharing, it does not by itself prevent privacy leakage through shared updates, nor does it ensure protection against malicious model tampering during the iterative aggregation process.
To address privacy leakage, differential privacy (DP) [9, 10] offers formal guarantees that the released outputs (e.g., clipped+noise-added gradients or differentially private final model) do not reveal sensitive information about any individual record. In practice, however, naively injecting noise can adversely impact accuracy in high-dimensional images, a phenomenon particularly problematic for reconstruction quality. Recent work explores refined mechanisms that adapt noise to internal model states to reduce utility loss [48, 49, 50]; yet these often introduce deployment overheads or rely on trust assumptions about the execution environment. Moreover, practical challenges — including floating-point implementation pitfalls [51, 52], precision-based attacks [53], and erroneous estimation of sensitivity [54]—can undermine DP guarantees if the DP solution is not properly designed. These challenges are amplified in image reconstruction tasks, where excessive noise can manifest as blur and loss of detail. Our proposed design is therefore directed toward the development of privacy mechanisms and training protocols that safeguard fine-grained FL-related structural information while ensuring stringent privacy guarantees.
In addition to privacy, we need appropriate integrity verification approaches to ensure that the global model has neither been replaced nor maliciously altered across training rounds. Watermarking and fingerprinting techniques [55] provide one way to authenticate the trained model and detect unauthorized modifications, allowing clients to verify the identity and authenticity of the downloaded global models before use. In higher-assurance settings, cryptographic protocols such as zero-knowledge proofs [56] can further attest to correct aggregation without exposing private updates. Together with FL-appropriate learning objectives (e.g., federated contrastive training [31]), these mechanisms can help establish a training loop in which neither raw data nor model parameters can be exploited for information leakage or manipulation, while still meeting the fidelity demands of complex reconstruction.
@@ II-B Federated Learning
III The Proposed PPFL-RDSN Framework
In this section, we overview our proposed PPFL-RDSN framework, illustrated in Figure 1:
  • Local Training: Within the PPFL-RDSN framework, each client carries out its local model training, maintaining exclusive access to its dataset without sharing it with other clients. During local training, Gaussian noise is strategically applied to the parameter updates to balance data utility and privacy. Additionally, each client embeds a watermark (one time) into its local RDSN model during training for subsequent integrity verification.
  • Model Aggregation and Distribution: After training, clients securely transmit their local models to a central server that acts as a secure aggregator. The server performs model aggregation using Federated Averaging (FedAvg) [32], combining local models into a global model based on weighted contributions.
  • Local EtLC: Once aggregated, the global model is distributed back to each client. Upon receipt, clients execute a model integrity check using their embedded watermark. After successful integrity verification, we input lossy or low-resolution images into the global model and carry out performance testing. Depending on the testing outcomes, the client either retains the global model or further initiates the next iteration of local training. Additionally, the Local EtLC subsystem allows clients to securely exchange encrypted and compressed low-resolution images. We assume pre-established shared secret keys for encrypted communication. Upon request, these images are decrypted, decompressed, and reconstructed into higher-resolution outputs using the verified global RDSN model. This process ensures collaborative enhancement while maintaining data privacy and integrity throughout the federated learning process.
III-A Threat Modeling and Mitigation
We analyze three classes of adversaries: 1) a honest-but-curious aggregator, 2) honest‑but‑curious clients (who may collude for inference while still following the protocol), and 3) a passive network eavesdropper. We focus the threat model on attacks that are countered by our current design.
Privacy Threats: Membership, inversion, and property inference attacks exploit shared gradients to uncover private training data [57, 58, 59]. These threats arise from both untrusted or semi-trusted aggregators as well as honest-but-curious clients that may pool their local observations and round-wise global models. Since our mechanism satisfies local $(\varepsilon,\delta)$-differential privacy, its guarantee is robust to any arbitrary side information—including potential collusion among clients or the aggregator because of the post-processing and composition properties of differential privacy [60, 11, 61].
Security Threats: Model substitution or parameter tampering attacks [62, 55] utilize the untrusted server to replace the global model with a modified tampered version; rollback or replays of old models form a related threat. Passive eavesdropping in transport can observe traffic, but cannot bypass end-to-end integrity checks; any active modification reduces to the substitution/tampering case. In our design, we do not claim our mechanism has the ability to defend against the following poisoning and backdoor attacks [63].